# Complexity Analysis of Normalizing Constant Estimation: from Jarzynski Equality to Annealed Importance Sampling and beyond

### v1

@@ E.4 An Upper Bound of the TV Distance along the OU Process
Assume that the target distribution $\pi\propto\mathrm{e}^{-V}$ satisfies Assump. 2, with the exception that $R\lesssim\frac{1}{\sqrt{\beta}}$. Let $\overline{\pi}_{\delta}$ be the distribution of $Y_{\delta}$ in the OU process (Eq. 3) initialized at $Y_{0}\sim\pi$, for some $\delta\lesssim 1$. Then,
$\operatorname{TV}(\pi,\overline{\pi}_{\delta})\lesssim\delta(\beta m^{2}+\beta Rm+d+\beta)+\delta^{\frac{1}{2}}d^{\frac{1}{2}}\beta(m+R).$
@@
Consider a simplified case where $R\ll 1$, $\beta\gtrsim 1$, and $m^{2}\asymp d$. Then it suffices to choose $\delta\lesssim\frac{\varepsilon^{2}}{\beta^{2}d^{2}}$ in order to guarantee $\operatorname{TV}(\pi,\overline{\pi}_{\delta})\lesssim\varepsilon$.
@@
Without loss of generality, suppose $\pi=\mathrm{e}^{-V}$. Let $\phi$ be the p.d.f. of $\operatorname{\mathcal{N}}\left(0,I\right)$, and define $\sigma^{2}:=1-\mathrm{e}^{-2\delta}\asymp\delta$. We will use the following inequality: $|\mathrm{e}^{a}-\mathrm{e}^{b}|\leq(\mathrm{e}^{a}+\mathrm{e}^{b})|a-b|$, which is due to the convexity of the exponential function. By the smoothness of $V$,
$\|\nabla V(x)\|=\|\nabla V(x)-\nabla V(x_{*})\|\leq\beta\|x-x_{*}\|\leq\beta(\|x\|+R).$
@@
$\lesssim\beta(\|x\|+R)\delta\|x\|+\beta\delta^{2}\|x\|^{2}$
$\lesssim\beta\delta\|x\|^{2}+\beta\delta R\|x\|.$
$\implies\operatorname{TV}(\pi,\pi^{\prime})$ $\lesssim\delta\int(\pi(x)+\pi^{\prime}(x))(\beta\|x\|^{2}+\beta R\|x\|+d)\mathrm{d}x.$
Note that
$\int\pi(x)(\beta\|x\|^{2}+\beta R\|x\|+d)\mathrm{d}x\leq\beta m^{2}+\beta Rm+d.$
Since $\operatorname{\mathbb{E}}_{\pi^{\prime}}\varphi=\operatorname{\mathbb{E}}_{\pi}\varphi(\mathrm{e}^{-\delta}\cdot)$, we also have $\int\pi^{\prime}(x)(\beta\|x\|^{2}+\beta R\|x\|+d)\mathrm{d}x\lesssim\beta m^{2}+\beta Rm+d$. We thus conclude that
$\operatorname{TV}(\pi,\pi^{\prime})\lesssim\delta(\beta m^{2}+\beta Rm+d).$
@@
$\lesssim\beta(\mathrm{e}^{\delta}\|x\|+R)\mathrm{e}^{\delta}\sigma\|u\|+\beta\mathrm{e}^{2\delta}\sigma^{2}\|u\|^{2}$
$\lesssim\beta(\|x\|+R)\delta^{\frac{1}{2}}\|u\|+\beta\delta\|u\|^{2}.$
@@
$\operatorname{TV}(\pi^{\prime},\overline{\pi}_{\delta})$ $\lesssim\beta\delta^{\frac{1}{2}}\iint(\pi^{\prime}(x+\sigma u)+\pi^{\prime}(x))(\|u\|\|x\|+\|u\|R+\delta^{\frac{1}{2}}\|u\|^{2})\phi(u)\mathrm{d}u\mathrm{d}x.$
Note that
$\iint\pi^{\prime}(x)(\|u\|\|x\|+\|u\|R+\delta^{\frac{1}{2}}\|u\|^{2})\phi(u)\mathrm{d}u\mathrm{d}x$
$\lesssim\operatorname{\mathbb{E}}_{\pi^{\prime}}\|\cdot\|d^{\frac{1}{2}}+Rd^{\frac{1}{2}}+\delta^{\frac{1}{2}}d\leq md^{\frac{1}{2}}+Rd^{\frac{1}{2}}+\delta^{\frac{1}{2}};$
$\iint\pi^{\prime}(x+\sigma u)(\|u\|\|x\|+\|u\|R+\delta^{\frac{1}{2}}\|u\|^{2})\phi(u)\mathrm{d}u\mathrm{d}x$
$=\iint\pi^{\prime}(y)(\|u\|\|y-\sigma u\|+\|u\|R+\delta^{\frac{1}{2}}\|u\|^{2})\phi(u)\mathrm{d}u\mathrm{d}y$
$\lesssim\iint\pi^{\prime}(y)(\|u\|\|y\|+\|u\|R+\delta^{\frac{1}{2}}\|u\|^{2})\phi(u)\mathrm{d}u\mathrm{d}y$
$\lesssim md^{\frac{1}{2}}+Rd^{\frac{1}{2}}+\delta^{\frac{1}{2}}.$
Therefore, $\operatorname{TV}(\pi^{\prime},\overline{\pi}_{\delta})\lesssim\beta\delta^{\frac{1}{2}}(md^{\frac{1}{2}}+Rd^{\frac{1}{2}}+\delta^{\frac{1}{2}})$. The proof is complete. ∎
@@ E.5 Discussion on the Overall Complexity of RDS
In RDS, an accurate score estimate $s_{\cdot}\approx\nabla\log\overline{\pi}_{\cdot}$ is critical for the algorithmic efficiency. Existing methods estimate scores through different ways. Here, we review the existing methods and their complexity guarantees for sampling, and leverage Thm. 11 to derive the complexity of normalizing constant estimation. Throughout this section, we always assume that the target distribution $\pi\propto\mathrm{e}^{-V}$ satisfies $m^{2}:=\operatorname{\mathbb{E}}_{\pi}\|\cdot\|^{2}<\infty$ and that $V$ is $\beta$-smooth.
@@
The seminal work directly leveraged the following Tweedie’s formula (Robbins, 1992) to estimate the score: Huang et al. (2024a)
@@
is the posterior distribution of $Y_{0}$ conditional on $Y_{t}=x$ in the OU process (Eq. 3). The paper proposed to sample from $\overline{\pi}_{0|t}(\cdot|x)$ by LMC and estimate the score via empirical mean, which has a provably better LSI constant than the target distribution $\pi$ (see Huang et al. (2024a, Lem. 2)). They show that if the scores $\nabla\log\overline{\pi}_{t}$ are uniformly $\beta$-Lipschitz, and assume that there exists some $c>0$ and $n>0$ such that for any $r>0$, $V+r\|\cdot\|^{2}$ is convex for $\|x\|\geq\frac{c}{r^{n}}$, then w.p. $\geq 1-\zeta$, the overall complexity for guaranteeing $\operatorname{KL}(\mathbb{Q}\|\mathbb{Q}^{\dagger})\lesssim\varepsilon^{2}$ is
@@
which is also the complexity of obtaining a $\widehat{Z}$ satisfying Eq. 5.
@@
Assume that $V$ is $\beta$-smooth, and the distributions along the OU process starting from $\pi\propto\mathrm{e}^{-V}$ and $\pi^{\prime}\propto\mathrm{e}^{-2V}$ have potentials whose Hessians are uniformly $\succeq\beta I$, then the complexity for guaranteeing $\operatorname{\mathbb{E}}\operatorname{KL}(\mathbb{Q}\|\mathbb{Q}^{\dagger})\lesssim\varepsilon^{2}$ is
@@ Appendix F Supplementary Lemmas
By applying Itô’s formula to Eq. 16 for $t\in[t_{k},t_{k+1}]$, we have
@@
Define $\widehat{\pi}_{\lambda}\propto\exp\left(-V-\frac{\lambda}{2}\|\cdot\|^{2}\right)$, $\lambda\geq 0$. Then under Assump. 2, $\operatorname{\mathbb{E}}_{\widehat{\pi}_{\lambda}}{\|\cdot\|^{2}}\leq m^{2}$ for all $\lambda\geq 0$.
@@
Lemma 34.
If a function $U$ on $\mathbb{R}^{d}$ satisfies $0\prec\nabla^{2}U\preceq\beta I$ for some $\beta>0$, and for any $t\geq 0$, let $x_{t}$ be the global minimizer of $U+\frac{t}{2}\|\cdot\|^{2}$. We have $\|x_{t}\|\leq\frac{\|x_{0}\|}{1+\frac{t}{\beta}}$.
Proof.
Since $\nabla U(x_{t})+tx_{t}=0$, taking time derivative yields $\nabla^{2}U(x_{t})\dot{x}_{t}+x_{t}+t\dot{x}_{t}=0$. Due to convexity, $\dot{x}_{t}=-(\nabla^{2}U(x_{t})+tI)^{-1}x_{t}$. Therefore,
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|x_{t}\|^{2}=x_{t}^{\mathrm{T}}\dot{x}_{t}=-x_{t}^{\mathrm{T}}(\nabla^{2}U(x_{t})+tI)^{-1}x_{t}\leq-\frac{\|x_{t}\|^{2}}{\beta+t},$
which implies $\frac{\mathrm{d}}{\mathrm{d}t}\left(\left(1+\frac{t}{\beta}\right)^{2}\|x_{t}\|^{2}\right)\leq 0$, and thus the proof is complete. ∎
Appendix G Review and Discussion on the Error Guarantee (Eq. 5)
@@ G.1 Literature Review of Existing Bounds
Traditionally, the statistical properties of an estimator are typically analyzed through its bias and variance. However, deriving closed-form expressions of the variance of $\widehat{Z}$ and $\widehat{F}$ in JE remains challenging. Recall that the estimator $\widehat{Z}=Z_{0}\mathrm{e}^{-W(X)}$, $X\sim\mathbb{P}^{\to}$ for $Z=Z_{0}\mathrm{e}^{-\Delta F}$, and that JE implies $\operatorname{Bias}\widehat{Z}=0$. For general (sub-optimally) controlled SDEs, Hartmann and Richter (2024) established both upper and lower bounds of the relative error of the importance sampling estimator, yet bounds tailored for JE are not well-studied. Inspired by this, we establish an upper bound on the normalized variance $\operatorname{Var}\frac{\widehat{Z}}{Z}$ in Prop. 35 at the end of this section using techniques in Rényi divergence. However, we remark that connecting this upper bound to the properties of the curve (e.g., action) is non-trivial, which we leave for future work.
@@ G.2 Equivalence in Complexities for Estimating $Z$ and $F$
We prove the claim in Rmk. 3 that estimating $Z$ with $O(\varepsilon)$ relative error and estimating $F$ with $O(\varepsilon)$ absolute error share the same complexity up to absolute constants. This follows directly from Lem. 27: for any $\varepsilon\in\left(0,\frac{1}{2}\right)$,
@@
G.3 Eq. 5 is Weaker than Bias and Variance
We demonstrate that Eq. 5 is a weaker criterion than controlling bias and variance, which is an immediate result from the Chebyshev inequality:
@@ G.4 An Upper Bound on the Normalized Variance of $\widehat{Z}$ in Jarzynski Equality
Under the setting of JE (Thm. 5), let $(v_{t})_{t\in[0,T]}$ be any vector field that generates $(\widetilde{\pi}_{t})_{t\in[0,T]}$, and define $\mathbb{P}$ as the path measure of Eq. 19. Then,
@@
We now bound $\operatorname{\mathbb{E}}_{\mathbb{P}}\left(\frac{\mathrm{d}\mathbb{P}^{\to}}{\mathrm{d}\mathbb{P}}\right)^{q}$ for any $q\in\mathbb{R}$. By Girsanov theorem (Lem. 1),
@@
where the last line is by the Cauchy-Schwarz inequality. Let $M_{t}:=-\sqrt{2}q\int_{0}^{t}\left\langle v_{r}(X_{r}),\mathrm{d}B_{r}\right\rangle$, $X\sim\mathbb{P}$ be a continuous martingale with quadratic variation $[M]_{t}=\int_{0}^{t}2q^{2}\|v_{r}(X_{r})\|^{2}\mathrm{d}r$. By Karatzas and Shreve (1991, Chap. 3.5.D), the process $t\mapsto\mathrm{e}^{M_{t}-\frac{1}{2}[M]_{t}}$ is a super martingale, and hence $\operatorname{\mathbb{E}}\mathrm{e}^{M_{T}-\frac{1}{2}[M]_{T}}\leq 1$. Thus, we have
@@
From Girsanov theorem (Lem. 15), we can similarly obtain the following RN derivative:
@@
and use the same argument to show that $\operatorname{\mathbb{E}}_{\mathbb{P}}\left(\frac{\mathrm{d}\mathbb{P}^{\leftarrow}}{\mathrm{d}\mathbb{P}}\right)^{q}$ has exactly the same upper bound as $\operatorname{\mathbb{E}}_{\mathbb{P}}\left(\frac{\mathrm{d}\mathbb{P}^{\to}}{\mathrm{d}\mathbb{P}}\right)^{q}$. In particular, we can use the same martingale argument, whereas now the backward continuous martingale is defined as $M^{\prime}_{t}:=-\sqrt{2}q\int_{t}^{T}\left\langle v_{r}(X_{r}),*\mathrm{d}B^{\leftarrow}_{r}\right\rangle$, $X\sim\mathbb{P}$, with quadratic variation $[M^{\prime}]_{t}=\int_{t}^{T}2q^{2}\|v_{r}(X_{r})\|^{2}\mathrm{d}r$. Therefore, we conclude that
@@
Appendix H Further Discussion on Related Works

### v2

@@ E.4 An Upper Bound of the TV Distance along the OU Process
Assume that the target distribution $\pi\propto\mathrm{e}^{-V}$ satisfies Assump.˜2. Let $\overline{\pi}_{\delta}$ be the distribution of $Y_{\delta}$ in the OU process ˜12 initialized at $Y_{0}\sim\pi$, for some $\delta\lesssim 1$. Then,
$\operatorname{TV}(\pi,\overline{\pi}_{\delta})\lesssim\delta(\beta m^{2}+d+\beta d)+\delta^{\frac{1}{2}}d^{\frac{1}{2}}\beta m.$
@@
Consider a simplified case where $\beta\gtrsim 1$ and $m^{2}\asymp d$. Then it suffices to choose $\delta\lesssim\frac{\varepsilon^{2}}{\beta^{2}d^{2}}$ in order to guarantee $\operatorname{TV}(\pi,\overline{\pi}_{\delta})\lesssim\varepsilon$.
@@
Without loss of generality, suppose $\pi=\mathrm{e}^{-V}$. Let $\phi$ be the p.d.f. of $\operatorname{\mathcal{N}}\left(0,I\right)$, and define $\sigma^{2}:=1-\mathrm{e}^{-2\delta}\asymp\delta$. We will use the following inequality: $|\mathrm{e}^{a}-\mathrm{e}^{b}|\leq(\mathrm{e}^{a}+\mathrm{e}^{b})|a-b|$, which is due to the convexity of the exponential function. By the smoothness of $V$, $\|\nabla V(x)\|=\|\nabla V(x)-\nabla V(0)\|\leq\beta\|x\|$.
@@
$\lesssim\beta\|x\|\delta\|x\|+\beta\delta^{2}\|x\|^{2}\lesssim\beta\delta\|x\|^{2}.$
$\implies\operatorname{TV}(\pi,\pi^{\prime})$ $\lesssim\delta\int(\pi(x)+\pi^{\prime}(x))(\beta\|x\|^{2}+d)\mathrm{d}x.$
Note that $\int\pi(x)(\beta\|x\|^{2}+d)\mathrm{d}x=\beta m^{2}+d$. Since $\operatorname{\mathbb{E}}_{\pi^{\prime}}\varphi=\operatorname{\mathbb{E}}_{\pi}\varphi(\mathrm{e}^{-\delta}\cdot)$, we also have
$\int\pi^{\prime}(x)(\beta\|x\|^{2}+d)\mathrm{d}x=\mathrm{e}^{-2\delta}\beta m^{2}+d\leq\beta m^{2}+d.$
We thus conclude that
$\operatorname{TV}(\pi,\pi^{\prime})\lesssim\delta(\beta m^{2}+d).$
@@
$\lesssim\beta\mathrm{e}^{\delta}\|x\|\mathrm{e}^{\delta}\sigma\|u\|+\beta\mathrm{e}^{2\delta}\sigma^{2}\|u\|^{2}$
$\lesssim\beta\|x\|\delta^{\frac{1}{2}}\|u\|+\beta\delta\|u\|^{2}.$
@@
$\operatorname{TV}(\pi^{\prime},\overline{\pi}_{\delta})$ $\lesssim\beta\delta^{\frac{1}{2}}\iint(\pi^{\prime}(x+\sigma u)+\pi^{\prime}(x))(\|u\|\|x\|+\delta^{\frac{1}{2}}\|u\|^{2})\phi(u)\mathrm{d}u\mathrm{d}x.$
Note that, first,
$\iint\pi^{\prime}(x)(\|u\|\|x\|+\delta^{\frac{1}{2}}\|u\|^{2})\phi(u)\mathrm{d}u\mathrm{d}x\lesssim\operatorname{\mathbb{E}}_{\pi^{\prime}}\|\cdot\|d^{\frac{1}{2}}+\delta^{\frac{1}{2}}d\leq md^{\frac{1}{2}}+\delta^{\frac{1}{2}}d;$
second,
$\iint\pi^{\prime}(x+\sigma u)(\|u\|\|x\|+\delta^{\frac{1}{2}}\|u\|^{2})\phi(u)\mathrm{d}u\mathrm{d}x$
$=\iint\pi^{\prime}(y)(\|u\|\|y-\sigma u\|+\delta^{\frac{1}{2}}\|u\|^{2})\phi(u)\mathrm{d}u\mathrm{d}y$
$\lesssim\iint\pi^{\prime}(y)(\|u\|\|y\|+\delta^{\frac{1}{2}}\|u\|^{2})\phi(u)\mathrm{d}u\mathrm{d}y\lesssim md^{\frac{1}{2}}+\delta^{\frac{1}{2}}d.$
Therefore, $\operatorname{TV}(\pi^{\prime},\overline{\pi}_{\delta})\lesssim\beta\delta^{\frac{1}{2}}d^{\frac{1}{2}}(m+\delta^{\frac{1}{2}}d^{\frac{1}{2}})$. The proof is complete. ∎
@@ E.5 Discussion on the Overall Complexity of RDS
In RDS, an accurate score estimate $s_{\cdot}\approx\nabla\log\overline{\pi}_{\cdot}$ is critical for the algorithmic efficiency. Existing methods estimate scores through different ways. Here, we review the existing methods and their complexity guarantees for sampling, and leverage Thm.˜5 to derive the complexity of normalizing constant estimation. Throughout this section, we always assume that the target distribution $\pi\propto\mathrm{e}^{-V}$ satisfies $m^{2}:=\operatorname{\mathbb{E}}_{\pi}\|\cdot\|^{2}<\infty$ and that $V$ is $\beta$-smooth.
@@
The seminal work directly leveraged the following Tweedie’s formula (Robbins, 1992) to estimate the score:
@@
is the posterior distribution of $Y_{0}$ conditional on $Y_{t}=x$ in the OU process ˜12. The paper proposed to sample from $\overline{\pi}_{0|t}(\cdot|x)$ by LMC and estimate the score via empirical mean, which has a provably better LSI constant than the target distribution $\pi$ (see Huang et al. (2024a, Lem. 2)). They show that if the scores $\nabla\log\overline{\pi}_{t}$ are uniformly $\beta$-Lipschitz, and assume that there exists some $c>0$ and $n>0$ such that for any $r>0$, $V+r\|\cdot\|^{2}$ is convex for $\|x\|\geq\frac{c}{r^{n}}$, then w.p. $\geq 1-\zeta$, the overall complexity for guaranteeing $\operatorname{KL}(\mathbb{Q}\|\mathbb{Q}^{\dagger})\lesssim\varepsilon^{2}$ is
@@
which is also the complexity of obtaining a $\widehat{Z}$ satisfying ˜1.
@@
Assume that $V$ is $\beta$-smooth, and the distributions along the OU process starting from $\pi\propto\mathrm{e}^{-V}$ and $\pi^{\prime}\propto\mathrm{e}^{-2V}$ have potentials whose Hessians are uniformly $\succeq-\beta I$, then the complexity for guaranteeing $\operatorname{\mathbb{E}}\operatorname{KL}(\mathbb{Q}\|\mathbb{Q}^{\dagger})\lesssim\varepsilon^{2}$ is
@@ Appendix F Supplementary Lemmas
By applying Itô’s formula to ˜14 for $t\in[t_{k},t_{k+1}]$, we have
@@
Define $\widehat{\pi}_{\lambda}\propto\exp\left(-V-\frac{\lambda}{2}\|\cdot\|^{2}\right)$, $\lambda\geq 0$. Then under Assump.˜2, $\operatorname{\mathbb{E}}_{\widehat{\pi}_{\lambda}}{\|\cdot\|^{2}}\leq m^{2}$ for all $\lambda\geq 0$.
@@
Appendix G Review and Discussion on the Error Guarantee ˜1
@@ G.1 Literature Review of Existing Bounds
Traditionally, the statistical properties of an estimator are typically analyzed through its bias and variance. However, deriving closed-form expressions of the variance of $\widehat{Z}$ and $\widehat{F}$ in JE remains challenging. Recall that the estimator $\widehat{Z}=Z_{0}\mathrm{e}^{-W(X)}$, $X\sim\mathbb{P}^{\to}$ for $Z=Z_{0}\mathrm{e}^{-\Delta F}$, and that JE implies $\operatorname{Bias}\widehat{Z}=0$. For general (sub-optimally) controlled SDEs, Hartmann and Richter (2024) established both upper and lower bounds of the relative error of the importance sampling estimator, yet bounds tailored for JE are not well-studied. Inspired by this, we establish an upper bound on the normalized variance $\operatorname{Var}\frac{\widehat{Z}}{Z}$ in Prop.˜3 at the end of this section using techniques in Rényi divergence. However, we remark that connecting this upper bound to the properties of the curve (e.g., action) is non-trivial, which we leave for future work.
@@ G.2 Equivalence in Complexities for Estimating $Z$ and $F$
We prove the claim in Rmk.˜1 that estimating $Z$ with $O(\varepsilon)$ relative error and estimating $F$ with $O(\varepsilon)$ absolute error share the same complexity up to absolute constants. This follows directly from Lem.˜9: for any $\varepsilon\in\left(0,\frac{1}{2}\right)$,
@@
G.3 ˜1 is Weaker than Bias and Variance
We demonstrate that ˜1 is a weaker criterion than controlling bias and variance, which is an immediate result from the Chebyshev inequality:
@@ G.4 An Upper Bound on the Normalized Variance of $\widehat{Z}$ in Jarzynski Equality
Under the setting of JE (Thm.˜1), let $(v_{t})_{t\in[0,T]}$ be any vector field that generates $(\widetilde{\pi}_{t})_{t\in[0,T]}$, and define $\mathbb{P}$ as the path measure of ˜17. Then,
@@
We now bound $\operatorname{\mathbb{E}}_{\mathbb{P}}\left(\frac{\mathrm{d}\mathbb{P}^{\to}}{\mathrm{d}\mathbb{P}}\right)^{q}$ for any $q\in\mathbb{R}$. By Girsanov theorem (Lem.˜2),
@@
where the last line is by the Cauchy-Schwarz inequality. Let $M_{t}:=-\sqrt{2}q\int_{0}^{t}\left\langle v_{r}(X_{r}),\mathrm{d}B_{r}\right\rangle$, $X\sim\mathbb{P}$ be a continuous martingale with quadratic variation $M]_{t}=\int_{0}^{t}2q^{2}\|v_{r}(X_{r})\|^{2}\mathrm{d}r$. By Karatzas and Shreve (1991, Chap. 3.5.D), the process $t\mapsto\mathrm{e}^{M_{t}-\frac{1}{2}[M]_{t}}$ is a super martingale, and hence $\operatorname{\mathbb{E}}\mathrm{e}^{M_{T}-\frac{1}{2}[M]_{T}}\leq 1$. Thus, we have
@@
From Girsanov theorem (Lem.˜4), we can similarly obtain the following RN derivative:
@@
and use the same argument to show that $\operatorname{\mathbb{E}}_{\mathbb{P}}\left(\frac{\mathrm{d}\mathbb{P}^{\leftarrow}}{\mathrm{d}\mathbb{P}}\right)^{q}$ has exactly the same upper bound as $\operatorname{\mathbb{E}}_{\mathbb{P}}\left(\frac{\mathrm{d}\mathbb{P}^{\to}}{\mathrm{d}\mathbb{P}}\right)^{q}$. In particular, we can use the same martingale argument, whereas now the backward continuous martingale is defined as $M^{\prime}_{t}:=-\sqrt{2}q\int_{t}^{T}\left\langle v_{r}(X_{r}),*\mathrm{d}B^{\leftarrow}_{r}\right\rangle$, $X\sim\mathbb{P}$, with quadratic variation $M^{\prime}]_{t}=\int_{t}^{T}2q^{2}\|v_{r}(X_{r})\|^{2}\mathrm{d}r$. Therefore, we conclude that
@@
Appendix H Related Works